### v1

# Dark Photon Dark Matter and Low-Frequency Gravitational Wave Detection with Gaia-like Astrometry

Abstract
Astrometric surveys offer us a method to search for elusive cosmic signatures, such as ultralight dark photon dark matter and gravitational waves, by observing the deflection to the apparent positions of the stars. The detection capabilities of such surveys rapidly decrease at low frequencies, because the signals become hardly distinguishable from the background motion of stars. In this work, we find that the background motion can be well described by a linear model over time, based on which we propose a linear background subtraction scheme. Compared to the conventional quadratic subtraction, the advantage of linear subtraction emerges within the frequency range below $6\times 10^{-9}~{}{\rm Hz}$. Taking dark photons with purely gravitational interactions, dark photons with additional $U(1)_{B}$ or $U(1)_{B-L}$ gauge interactions, and low-frequency gravitational waves as examples, we illustrate that the linear subtraction scheme can result in an enhancement of more than one order of magnitude in the exclusion limits of Gaia-like experiments in the low-frequency range.
preprint: DESY-24-106
I Introduction
The development of astrophysical observation methods has revolutionized our understanding of the cosmos, enabling us to probe subtle effects induced by new astrophysical sources or new physics with unprecedented precision. Recently, several Pulsar Timing Array (PTA) collaborations reported first evidence for the stochastic gravitational wave (GW) background at nano-Hertz (nHz) frequencies  Allen (1996); Agazie et al. (2023a); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023), which may arise from a population of inspiraling supermassive black hole binaries Sesana et al. (2008).
Apart from the GW signals, PTA experiments can also probe several new physics models. One example would be the ultralight bosonic dark matter Schive et al. (2014); Hui et al. (2017) which is capable of solving the so called small-scale problems Weinberg et al. (2015) of the cold dark matter model. Ultralight dark photon dark matter (DPDM) assumes that dark matter is composed of vector fields Adshead and Lozanov (2021); Amin et al. (2022) with a mass around $10^{-22}~{}{\rm eV}$. Due to its wave nature, the field value of the ultralight DPDM oscillates in an approximate coherent manner. Therefore, the ultralight DPDM can induce nontrivial oscillations of the metric due to the pure gravitational effects Khmelnitsky and Rubakov (2014); Nomura et al. (2020); Chen et al. (2023), or impose external oscillating forces on the test objects via additional U(1) gauge interactions Pierce et al. (2018), both of which can modify the pulse arrival time Porayko et al. (2018); Xue et al. (2022); Afzal et al. (2023); Nelson and Scholtz (2011); Wu et al. (2022); Xia et al. (2023).
On the other hand, the GW or DPDM background can result in the deflection of light emitted from the astrophysical objects to the Earth, due to the following two distinct effects:
  • GWs or DPDM with purely gravitational interactions can induce metric perturbations, which can affect the trajectory of light emitted from the sources Braginsky et al. (1990); Pyne et al. (1996); Kaiser and Jaffe (1997); Kopeikin et al. (1999); Book and Flanagan (2011). As a result, the apparent positions of stars we observe slightly deviate from their initial positions. Such displacements depend on metric perturbations at both the Earth and stars, while the Earth term should be dominant, as we will explain later. It should also be mentioned that, unlike the deflection induced by ultralight scalar dark matter, which is suppressed due to an approximate spherical symmetry, the deflection caused by DPDM is generally not suppressed.
  • DPDM coupling to the baryon number ${\rm U}(1)_{B}$ or the baryon number minus lepton number ${\rm U}(1)_{B-L}$ can lead to the aberration effect, which also changes the apparent position of stars Guo et al. (2019).
By precisely measuring the variation of stars’ apparent positions, it is hopeful to capture the traces of GW and DPDM signals in the Gaia mission Prusti et al. (2016) and upcoming astrometric surveys like Roman Wang et al. (2021, 2022); Haiman et al. (2023); Pardo et al. (2023) and Theia Boehm et al. (2017); Malbet et al. (2021); Garcia-Bellido et al. (2021). Gaia observes positions, proper motions, and parallaxes of more than $10^{9}$ objects with unprecedented accuracy Prusti et al. (2016). The sensitive frequency of Gaia is determined by its survey lifetime $T_{\rm obs}=10~{}{\rm year}$ and observational cadence $\Delta t=24~{}{\rm day}$, typically, $10^{-9}~{}{\rm Hz}\lesssim f\lesssim 10^{-7}~{}{\rm Hz}$. Hence Gaia can provide complementary detection capabilities to PTA experiments for nHz astrophysical sources Moore et al. (2017); Klioner (2018); Qin et al. (2019); O’Beirne and Cornish (2018); Mihaylov et al. (2018); Bini and Geralico (2018); Darling et al. (2018); Mihaylov et al. (2020); Guo et al. (2019); Qin et al. (2021); Aoyama et al. (2021); Jaraba et al. (2023); Liang et al. (2024); Çalışkan et al. (2024).
To extract the deflections to the apparent positions of stars caused by external signals, it is crucial to subtract their intrinsic background motion stemming from the physical motion of stars with respect to the solar barycenter and the secular aberration due to the moving reference system. This intrinsic motion can induce large angular deflection far exceeding the resolution of Gaia. For the signals with frequencies $f\gtrsim 2/T_{\rm obs}\sim 6\times 10^{-9}~{}{\rm Hz}$, the background subtraction essentially does not influence the sensitivity, since the oscillation pattern of GWs or DPDM is distinctive from the background. Conversely, if $f\lesssim 6\times 10^{-9}~{}{\rm Hz}$, a half oscillating cycle would not be observable in the time-series data within the operational lifetime of Gaia. Then the signal is also partially subtracted, thereby reducing the sensitivity of detecting low-frequency GW and DPDM signals. Analogous to PTAs, the quadratic model is commonly used to subtract the background proper motion in the existing literature Moore et al. (2017); Guo et al. (2019); Wang et al. (2021, 2022). However, a more comprehensive study of the background motion of stars and subtraction schemes should be implemented.
In this work, we propose a linear background subtraction scheme for the time-series data of apparent stellar positions in Gaia-like astrometric surveys. Upon analyzing the background motion of stars, we find that the impact on the apparent stellar position due to the constant acceleration from the galactic gravitational potential is negligibly small relative to Gaia’s sensitivity. Meanwhile, the galactocentric acceleration of the solar barycenter mimics a linear term over time in the proper motion due to aberration. These observations serve as the primary basis for our linear subtraction approach. Nevertheless, unresolvable binary systems could bring additional noise, which cannot be effectively subtracted. We provide concrete examples by calculating the astrometric deflection induced by DPDM with purely gravitational interactions, while also revisiting scenarios involving DPDM with ${\rm U}(1)_{B}$ or ${\rm U}(1)_{B-L}$ gauge symmetry and GWs from binary systems. Subsequent numerical simulation using mock data reveals that our linear subtraction scheme significantly enhances Gaia’s detection capabilities in the low-frequency range, improving sensitivity by more than an order of magnitude compared to the quadratic subtraction.
The layout of this paper is as follows. In Sec. II we illustrate the background motion of stars can be well described by a linear model over time. In Sec. III we calculate astrometric deflections induced by GWs and DPDM. The numerical simulation and results are presented in Sec. IV. We summarize our main conclusions in Sec. V.
II Background motion of the apparent stellar locations
Astrometric objects have background motion around the galactic center. Their motion is randomly distributed with a potentially non-zero expectation value depending on the star’s sky location, due to the common motion of the galactic disk. We assume that the velocity $\boldsymbol{v}=(v_{r},v_{\theta},v_{\phi})$ in spherical polar coordinates at the galactic center follows a Gaussian distribution
$P(\boldsymbol{v}|\boldsymbol{n})=\prod_{i=1}^{3}\frac{1}{\sqrt{2\pi}\sigma_{v,i}}\exp\left(\frac{(v_{i}-v_{0,i}(\boldsymbol{n}))^{2}}{2\sigma_{v,i}^{2}}\right)\;,$ (1)
where $\boldsymbol{n}$ denotes the sky location of stars, $v_{0,i}(\boldsymbol{n})$ represents the expectation value of $\boldsymbol{v}$ in each direction, and $\sigma_{v,i}$ refers to the velocity dispersion. The anisotropy of the velocity dispersion has been observed by SDSS and Gaia. Approximately we have the relation $\sigma_{v,r}\simeq 2\sigma_{v,\theta/\phi}$ for stars with high metallicity, which may originates from a major merger event of a satellite galaxy between $8$ and $11$ Gyr ago Belokurov et al. (2018). For qualitative discussions, we assume the uncertainty $\sigma_{v}\simeq 100\,{\rm km}/{\rm s}\simeq 1.01\times 10^{-4}\,{\rm pc/year}$ for all stars on both directions on the sky sphere. Choosing a typical distance of $1$ kpc, [^1] The peak distribution of distances between the stars and us observed by Gaia is $1\sim 2$ kpc Vallenari et al. (2023). the angular deflection observed over $10$ years reaches $\simeq 1.01\times 10^{-6}\,{\rm rad}$, which significantly exceeds the single-exposure resolution $\sigma_{\rm ss}\simeq 100\,\mu{\rm as}=4.85\times 10^{-10}\,{\rm rad}$ of Gaia for a single star. It then becomes necessary to subtract the background proper motion from the data before searching for the GW or DPDM signal.
Let us consider stars in two categories, isolated stars and the ones in binary or higher-order systems. For isolated stars, their velocities in general change over time. Hence we should also estimate their accelerations, which mainly include contributions from two aspects:
  • The constant accelerations of stars. The density gradient in the Milky Way could induce constant accelerations for stars. Unlike their proper motion, the constant accelerations of stars are predictable due to the galactic mass profile, indicating the uncertainty $\sigma_{a}$ should be much smaller than the expectation values $|\boldsymbol{a}_{0}|$. Typical stars in the Milky Way experience an acceleration of $|\boldsymbol{a}_{0}|\sim 10^{-10}\,{\rm m}/{\rm s}^{2}$ Silverwood and Easther (2019), which leads to a displacement of $|\boldsymbol{a}_{0}|\Delta t^{2}/2\simeq 1.6\times 10^{-10}$ pc over the 10-year observation. For a star located 1 kpc away from us, the corresponding angular deflection is about $0.03\,\mu{\rm as}$, significantly below the detection threshold.
  • The constant acceleration of solar system barycenter. Like other stellar systems, the velocity of our solar system is not precisely constant. Instead, it follows a curved orbit in the Galaxy. For the observer, the nonzero acceleration of the solar system barycenter gives rise to the aberration effect, which can be considered as extra (spurious) proper motion superimposed on stars Kopeikin and Makarov (2006); Klioner et al. (2021)
    $\frac{{\rm d}\delta\boldsymbol{n}}{{\rm d}t}=\boldsymbol{a}_{\rm SSB}-(\boldsymbol{a}_{\rm SSB}\cdot\boldsymbol{n})\boldsymbol{n}\;,$ (2)
    where $\delta\boldsymbol{n}$ denotes the displacement in the apparent stellar position due to aberration, and $\boldsymbol{a}_{\rm SSB}$ is the acceleration of the solar system barycenter (SSB). It then becomes evident that $\boldsymbol{a}_{B}$ only introduces a modification that is linearly dependent on time.
In short, the background motion of the apparent stellar locations is governed by the proper motion of stars, which can be significantly greater than the resolution of Gaia. The acceleration of stars is very minor and thus plays a negligible role. Moreover, the acceleration of the solar system barycenter yields spurious proper motion. The above backgrounds can be eliminated through a linear subtraction scheme.
On the other hand, the unresolved binary or higher-order systems in the Galaxy could produce a significant background motion for the relevant stars. In the practical numerical calculation, we employ a data compression technique akin to the one described in Refs. Moore et al. (2017); Klioner (2018), since the vast number of objects ($\sim 10^{9}$) in the full dataset of Gaia severely slows down the process of numerical simulation. Specifically, we compress the data of $10^{5}$ stars in the same sky area into one virtual star. For each star, the probability of being part of a binary or high-order system and having specific orbital parameters is statistically independent of the status of the other stars. After data compression, the astrometric data of $10^{5}$ stars in each sky area are added together. The background motion of individual stars in binary or higher-order systems manifests as a time-dependent noise of the compressed virtual star.
A survey of nearby ($<250$ pc) stellar objects finds $\sim 55\%$ solar-type stars are isolated, while the remaining $\sim 45\%$ stars are in binary or higher-order systems Raghavan et al. (2010). Here we take the binary systems for simplicity. Following the Kepler’s law, the orbital period $P_{b}$ has the following relation with the semi-major axis $a_{b}$
$a_{b}=\left[\frac{Gm_{1}(1+q)P_{b}^{2}}{(2\pi)^{2}}\right]^{1/3}\;,$ (3)
where $m_{1}$ labels the mass of the primary star and $q=m_{2}/m_{1}<1$. If both stars are observed by Gaia, the neat contribution from the binary to the displacement of stars will approximate to
$r_{\rm neat}\simeq a_{b}(1+e_{b})\frac{1-q}{1+q}\;,$ (4)
where $e_{b}$ is the orbit eccentricity. Then the neat contribution will be fully canceled if $q=1$.
Figure 1: The result of 10000 times simulation of the injected white noise (WN), proper motion (PM) and binary motion. Dashed lines correspond to the cases with white noise removed. The input Gaia operational parameters align with the discussion in Sec. IV.
For the $q<1$ case, we illustrate by considering a $e_{b}=0$ binary composed of two stars with the total mass $M=m_{1}+m_{2}$. The neat amplitude of the observed angular deflection in one direction is
$\begin{split}A_{b}=\frac{r_{\rm neat}\lambda(\iota_{b},\psi_{b})}{\sqrt{2}D}&=6.10\times 10^{-9}{\rm rad}\left(\frac{1-q}{1+q}\right)\frac{\lambda(\iota_{b},\psi_{b})}{\sqrt{2}}\\ &\times\left[\frac{P_{b}}{\rm year}\right]^{2/3}\left[\frac{D}{1{\rm kpc}}\right]^{-1}\left[\frac{M}{2M_{\odot}}\right]^{1/3}\;,\end{split}$ (5)
where $D$ is the distance between the observer and the binary barycenter. Also, $\lambda(\iota_{b},\psi_{b})=\cos\iota_{b}\sin\psi_{b}+\cos\psi_{b}$ with $\iota_{b}$ and $\psi_{b}$ being the inclination angle and the polarization angle of the stellar binary plane, respectively. Eq. (5) indicates binaries with $P_{b}\gtrsim 1\,{\rm year}$ and $D\sim 1\,{\rm kpc}$ may be resolved by Gaia. According to Raghavan et al. (2010); Simonetti et al. (2020), the period distribution approximately follows the log-Normal distribution
$P(\log_{10}P_{b}/{\rm day})\simeq\frac{1}{2.28\sqrt{2\pi}}e^{-\frac{\left(\log_{10}P_{b}/{\rm day}-5.03\right)^{2}}{2\times 2.28^{2}}}\;,$ (6)
which centers around a period of $293.6$ years. About $20\%$ stellar binaries have the orbital periods from $24$ days to $10$ years, which would induce an irreducible background for the $10$-year observation of Gaia mission.
We calculate the power spectral density (PSD) originating from binary stars $S_{\rm bin}(f)$ by implementing a Monte-Carlo simulation, where we assume $45\%$ of stars are in binaries. We use Eq. (5) with $D=1$ kpc, while the orbital period is drawn from Eq. (6). Further, $\cos\iota_{b}$, $\psi_{b}$ and $q$ are treated as free parameters, randomly selected from uniform distributions within the intervals of $[-1,1)$, $[0,2\pi)$ and $[0.1,1)$, respectively. In each simulation, we generate astrometric data of $10^{5}$ stars, which are then compressed into one virtual star. We calculate the PSD of astrometric data of the virtual star. The final $S_{\rm bin}(f)$ is obtained by averaging the PSD over all simulations. The total noise PSD $S_{n}(f)$ is the combination of $S_{\rm bin}(f)$ and that from the instrumental noise $S_{w}(f)$, namely,
$S_{n}(f)=S_{w}(f)+S_{\rm bin}(f)\;.$ (7)
Our result is shown in Fig. 1, where one can observe that the linear subtraction approach effectively reduces the noise from stellar proper motion, reproducing the PSD of pure white noise (WN) above $10^{-8}$ Hz. However, for lower frequency, part of the WN is also removed by the linear subtraction, leading to a smaller PSD. Furthermore, the inclusion of binaries indeed introduces significant noise when $f\lesssim 10^{-7}\,{\rm Hz}$, corresponding to $P_{b}\gtrsim 24$ day. This noise grows as $P_{b}$ increases, coinciding with the behavior manifested by Eq. (5). As $f$ falls below $6\times 10^{-9}\,{\rm Hz}$, the observed angular deflection $A_{b}$ of the binary system no longer exhibits periodic oscillations during the duration of Gaia’s observations. Consequently, two subtraction schemes will reduce the time-dependent noise induced by unresolved binaries in the low-frequency region, with the quadratic subtraction being more efficient. However, the quadratic subtraction also leads to more reductions to our signal at low frequencies, and we will see that the linear subtraction achieves better signal-to-noise ratio eventually.
III Astrometric deflections induced by GW and DPDM
III.1 GWs from binary systems
To begin with, we first revisit the formula for light deflection induced by GWs Book and Flanagan (2011); Pyne et al. (1996). The presence of GWs triggers metric perturbations, which lead to slight deflection of light traveling from the star to the Earth, ultimately resulting in the apparent displacement of stellar positions. By repeatedly measuring the apparent positions of a large number of stars across the sky, Gaia might be capable of capturing such deflection and thus extracting the characteristic patterns of GWs.
The variation in the line of sight direction $n^{i}$ due to GWs can be expressed as Pyne et al. (1996); Book and Flanagan (2011)
$\delta n^{i}\simeq\frac{n^{i}-p^{i}}{2(1-\bm{n}\cdot\bm{p})}h_{jk}({\rm E})n^{j}n^{k}-\frac{1}{2}h_{ij}({\rm E})n^{j}\;,$ (8)
where the unit vector $\bm{p}$ described by two rotational angles $\theta_{p}$ and $\varphi_{p}$ represents the propagated direction with $p^{i}$ being its components, and $h_{ij}(\rm E)$ refers to the metric perturbation calculated at the Earth. The metric perturbation $h_{ij}$ caused by monochromatic GWs with frequency $f_{\rm GW}$ reads
$h_{ij}(t,\boldsymbol{x})={\rm Re}\left[\mathcal{H}^{\rm GW}_{ij}e^{2\pi if_{\rm GW}(t-{\bm{p}}\cdot\bm{x})}\right]\;,$ (9)
with $\mathcal{H}^{\rm GW}_{ij}$ denoting the amplitude tensor. The GW possesses two polarization directions, and thus $\mathcal{H}^{\rm GW}_{ij}$ can be written as the following form
$\mathcal{H}^{\rm GW}_{ij}=h_{+}\epsilon^{+}_{ij}(\boldsymbol{p},\psi)e^{i\phi_{+}}+h_{\times}\epsilon^{\times}_{ij}(\boldsymbol{p},\psi)e^{i\phi_{\times}}\;,$ (10)
where $\epsilon^{+}_{ij}$ and $\epsilon^{\times}_{ij}$ are two polarization tensors depending on $\boldsymbol{p}$ and the GW polarization angle $\psi$ Arzoumanian et al. (2023), $h_{+}$ and $h_{\times}$ are their amplitudes, and $\phi_{+}$ and $\phi_{\times}$ are two corresponding phases. We consider the monochromatic GW radiated from a binary system with a circular orbit, the amplitudes $h_{+}$ and $h_{\times}$ of which can be respectively expressed as
$h_{+}=h^{\rm GW}_{0}\frac{1+\cos^{2}\iota}{2}\;,\quad h_{\times}=h^{\rm GW}_{0}\cos\iota\;,$ (11)
where $h^{\rm GW}_{0}$ is a constant depending on the chirp mass, the angular orbital frequency and the luminosity distance to the source, and $\iota$ is the inclination angle of the binary. Finally, for the circular orbit, the relation $\varphi_{+}-\varphi_{\times}=\pi/2$ should be satisfied.
We emphasize that we adopt the distant source approximation, thus we only need to keep the “Earth term” in the metric perturbation in Eq. (8). This approximation is valid when the sources are located at distances much larger than the wavelength of GWs Book and Flanagan (2011); Pyne et al. (1996); Moore et al. (2017). Under this approximation, the perturbations at different sources are uncorrelated, resembling a kind of random noise. Such noise is much smaller than the Gaia’s resolution and can be safely omitted.
III.2 DPDM with purely gravitational interactions
Analogous to GWs, dark photon fields can also induce metric fluctuations via gravitational effects Nelson and Scholtz (2011); Khmelnitsky and Rubakov (2014); Nomura et al. (2020). In this subsection, we derive the angular deflection of light emitted from the star as it passes through DPDM with purely gravitational interactions due to metric perturbations Çalışkan et al. (2024). We should mention that DPDM with a mass below $10^{-22}$ eV is subject to severe constraints from, e.g., the integrated Sachs-Wolfe effect on CMB anisotropies ($\gtrsim 10^{-24}~{}{\rm eV}$Hlozek et al. (2015), the stellar kinematics of dwarf spheroidal galaxies (a few times of $10^{-23}$ eV) González-Morales et al. (2017); Kendall and Easther (2020) and the Lyman-$\alpha$ forest ($\gtrsim 10^{-21}~{}{\rm eV}$Hui et al. (2017); Armengaud et al. (2017). Though the lower bound on ultralight dark matter mass is still under debate, here we assume that dark photons with $m_{A}\lesssim 10^{-22}~{}{\rm eV}$ constitute at most 10% of the total dark matter energy density.
We first describe the DPDM model. The dark photon field is expressed as
${\bm{A}}(t,{\bm{x}})\simeq A_{0}(\boldsymbol{x}){\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})\exp(im_{A}t)\right]\;,$ (12)
where the polarization vector $\boldsymbol{q}(\boldsymbol{x})=(\cos\theta_{q}\cos\varphi_{q}e^{i\phi_{1}}$, $\cos\theta_{q}\sin\varphi_{q}e^{i\phi_{2}},\sin\theta_{q}e^{i\phi_{3}})^{\rm T}$ is parametrized by two polarization angles $\theta_{q}$ and $\varphi_{q}$ and three phases $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$, all of which are in general certain functions of $\boldsymbol{x}$. Under the synchronous gauge, the metric perturbation has the following form
$h_{ij}^{\rm osc}(t,\boldsymbol{x})\simeq{\rm Re}\left[\mathcal{H}_{ij}^{\rm DP}(\boldsymbol{x})e^{2im_{A}t}\right],$ (13)
where
$\mathcal{H}_{ij}^{\rm DP}(\boldsymbol{x})=-h^{\rm osc}_{0}\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right].$ (14)
The signal amplitude can be written as $h^{\rm osc}_{0}=4\pi GA^{2}_{0}(\boldsymbol{x})$, with $G$ and $A_{0}(\boldsymbol{x})$ being respectively the gravitational constant and the amplitude of dark photon field. The amplitude of DPDM is
$A_{0}(\boldsymbol{x})=\frac{\sqrt{2\rho_{\rm DM}(\boldsymbol{x})}}{m_{A}}\;,$ (15)
where $\rho_{\rm DM}(\boldsymbol{x})$ is the energy density of the DPDM. A detailed derivation of Eq. (13) can be found in appendix A. As a result, we can estimate the signal amplitude
$h^{\rm osc}_{0}=5.2\times 10^{-15}\left(\frac{\rho_{\rm DM}}{0.4\,{\rm GeV}/{\rm cm}^{3}}\right)\left(\frac{10^{-23}\,{\rm eV}}{m_{A}}\right)^{2}.$ (16)
It is known that the metric perturbation $h_{ij}$ given in Eq. (13) yields a delay of the arrival time of the pulsar light, which can be probed by PTAs Nomura et al. (2020); Xue et al. (2022); Wu et al. (2022); Xia et al. (2023). Here we show that the same metric perturbation could also lead to the deflection of star’s proper position, namely,
$\delta n^{i}\simeq-\frac{(\delta^{ik}-n^{i}n^{k})}{2}n^{j}h_{jk}({\rm E})+\mathcal{O}(v_{\rm vir})\;,$ (17)
where the perturbation $h_{jk}({\rm E})$ is again calculated at the Earth. One can obtain the above expression by taking the limit $\boldsymbol{p}\sim v_{\rm vir}\ll 1$ in Eq. (8). The complete gauge-invariant expressions of $\delta n^{i}$ can be found in appendix B. In this case the star terms are suppressed by the distance and the virial velocity, as can be seen from Eqs. (71) and (76).
III.3 DPDM with additional gauge couplings
DPDM can also interact with ordinary matter if they both carry ${\rm U(1)}_{B}$ or ${\rm U(1)}_{B-L}$ charge. As a consequence, any object could experience an oscillating force due to the DPDM background Pierce et al. (2018)
$\bm{a}(t,\bm{x})\simeq\epsilon e\frac{q_{\rm obj}}{m_{\rm obj}}\frac{\partial}{\partial t}\bm{A}(t,\boldsymbol{x})\;,$ (18)
where $\epsilon$ is the coupling constant of the dark photon, $e$ denotes the electromagnetic coupling constant, and $q_{\rm obj}$ refers to $B$ or $B-L$ charge the object carries, and $m_{\rm obj}$ is the mass of the test object.
The above acceleration leads to a velocity variation of the object
$\delta{\bm{v}}(t,{\bm{x}})=\epsilon e\frac{q_{\rm obj}}{m_{\rm obj}}\bm{A}(t,\boldsymbol{x})\;,$ (19)
which can cause aberrations of the stellar proper motion due to the moving observer Guo et al. (2019). The resulting apparent angular deflection can be expressed as
$\delta\boldsymbol{n}(t,\boldsymbol{n})=\delta\boldsymbol{v}(t,\boldsymbol{x}_{\rm E})-(\delta\boldsymbol{v}(t,\boldsymbol{x}_{\rm E})\cdot\boldsymbol{n})\boldsymbol{n}\;.$ (20)
where $\boldsymbol{x}_{\rm E}$ is the coordinate of the observer. Moreover, the dark photon field can also directly lead to stellar motion, with an amplitude approximated by $\delta\boldsymbol{n}\lesssim|\delta\boldsymbol{v}|\cdot t/D$. However, for the typical observation time $(\sim 1~{}{\rm year})$ and distance $D\gtrsim 1~{}{\rm kpc}$, $\delta\boldsymbol{n}$ is rather tiny and will not be taken into consideration Guo et al. (2019).
IV Sensitivity of GW and DPDM searches with Gaia
In this section, we estimate the sensitivity of Gaia-like astrometry for GW and DPDM induced signals. We consider both Gaia and its next-generation upgrade, referred to as “XG-Gaia”. The operational parameters of these missions are set as follows:
$\begin{split}{\rm Gaia}:&\quad\sigma_{\rm ss}=100~{}\mu{\rm as}\;,~{}\Delta t=24~{}{\rm day}\;,\\ &\quad T_{\rm obs}=10~{}{\rm year}\;,N=1\times 10^{9}\;;\\ \text{XG-Gaia}:&\quad\sigma_{\rm ss}=1~{}\mu{\rm as}\;,~{}\Delta t=24~{}{\rm day}\;,\\ &\quad T_{\rm obs}=20~{}{\rm year}\;,N=1\times 10^{9}\;,\end{split}$
where $\sigma_{\rm ss}$ is the noise from a single exposure, [^2] The uncertainties on the proper motion parameters depend on stellar magnitudes $G$. It is expected the final uncertainty $\sigma_{\rm ss}\simeq 100~{}\mu{\rm as}$ for stars with $G\simeq 20$ in Gaia DR5 Vallenari et al. (2023). As these stars constitute the vast majority of the Gaia dataset, we will use $\sigma_{\rm ss}=100~{}\mu{\rm as}$ for all stars later on. $\Delta t$ is the observational cadence, $T_{\rm obs}$ is the total observation time, and $N$ is the total number of stars considered in the analysis.
As previously mentioned, we compress the data of $10^{5}$ stars in the same sky area by randomly generating $\widetilde{N}=10^{4}$ virtual stars uniformly distributed across the sky. Consequently, the instrument error for the measurement on each virtual star is rescaled to $\widetilde{\sigma}=\sigma_{\rm ss}/\sqrt{10^{9}/10^{4}}$. To estimate the sensitivity in detecting GWs and DPDM with Gaia-like experiments, we construct a mock dataset by injecting a specific GW or DPDM signal into the virtual stars. In all cases, it has been shown that the signals we considered are always dominated by the Earth term. Therefore, the astrometric effects on the stars in one area, caused by GWs or DPDM, can be approximated to the global effect on the virtual stars within that area.
In addition to the signal component, in Sec. II we have established that the background motion due to stellar proper motion and the acceleration of the SSB is effectively removed by linear subtraction. The constant acceleration resulting from the galaxy’s mass distribution is sufficiently small, thus can be safely ignored. Since linear subtraction is consistently applied to the astrometric data of virtual stars, background motion is not injected into the analysis. To address the time-dependent noise caused by stellar binaries, we adopt two different strategies for Gaia and XG-Gaia:
Figure 2: The 95% exclusion limits from Gaia (green lines) and XG-Gaia (red lines) on the GW strain $h_{0}$ from a binary system, the energy density $\rho_{\rm DM}$ of the DPDM with purely gravitational interactions, and the coupling constant $\epsilon$ of $U(1)_{B}/U(1)_{B-L}$ DPDM, respectively. Model parameters regarding all the angles have been integrated out. Solid lines represent the bounds with linear subtraction, while dashed lines correspond to quadratic subtraction. In addition, we use blue lines to denote the results including the time-dependent noise from unresolved binary systems. For comparison, constraints from NANOGrav 15-year Agazie et al. (2023b); Afzal et al. (2023), EPTA DR2 Smarra et al. (2023), and PPTA DR2 Xue et al. (2022); Wu et al. (2022) results are labeled by orange, cyan and purple curves, respectively, with the solid (dashed) curves being the (un)correlated constraints. In the top-right panel, the gray dashed line shows the amount of local dark matter density near the Earth as a reference. In the bottom two panels, the horizontal solid lines are the limits from the MICROSCOPE WEP experiment Touboul et al. (2017, 2022); Amaral et al. (2024).
  • For Gaia, we assume that all stellar binaries are unresolvable. Thus we use the PSD $S_{\rm bin}(f)$ calculated in Sec. II, as shown in Fig. 1.
  • For XG-Gaia, with $\sigma_{\rm ss}=1~{}\mu{\rm as}$, binaries with net amplitudes $A_{b}\lesssim\sigma_{\rm ss}$ constitute only 0.01% according to Eqs. (5) and (6). Therefore, we assume that almost all binaries have been resolved and removed from the data. Consequently, only single stars ($\sim$55of the total stars) are used for numerical simulation, and $S_{\rm bin}(f)=0$.
We construct $\chi^{2}$-functions for both the signal hypothesis and null hypothesis in the frequency domain, namely,
$\chi^{2}_{s}=\frac{2}{T_{\rm obs}}\sum_{I,K}\frac{|\widetilde{\bm{d}}_{I}(f_{K})-\widetilde{\bm{s}}_{I}(\Phi;f_{K})|^{2}}{S_{n}(f_{K})}\;,$ (21)
and
$\chi^{2}_{n}=\frac{2}{T_{\rm obs}}\sum_{I,K}\frac{|\widetilde{\bm{d}}_{I}(f_{K})|^{2}}{S_{n}(f_{K})}\;,$ (22)
where $\widetilde{\bm{d}}_{I}$ is the astrometry data of the virtual star $I$ after linear or quadratic subtraction, $\widetilde{\bm{s}}_{I}(\Phi;f_{K})$ represents the model predictions for a given set of parameters $\Phi$ after the same subtraction strategy, and $S_{n}(f)$ is defined in Eq. (7) with $S_{w}(f)=\widetilde{\sigma}\Delta t$. Note that both $\widetilde{\bm{d}}_{I}$ and $\widetilde{\bm{s}}_{I}$ are the quantities after Fourier transforms, and we sum over $I$ and $K$, which correspond to the number of virtual stars and that of bins in the PSD, respectively. The difference between the signal hypothesis and null hypothesis, i.e., $\Delta\chi^{2}=\chi^{2}_{s}-\chi^{2}_{n}$ becomes an indicator for measuring the significance of the signal.
The free parameters in $\Phi$ for (A) GWs from binary systems, (B) DPDM with purely gravitational interactions, and (C) DPDM with ${\rm U}(1)_{B}$ or ${\rm U}(1)_{B-L}$ gauge symmetry are respectively
$\begin{split}\text{(A)}:&\quad\{f_{\rm GW},\;h_{0},\;\iota,\;\phi_{+},\;\theta_{p}\;,\varphi_{p},\;\psi\}\;;\\ \text{(B)}:&\quad\{m_{A},\;\rho_{\rm DM},\;\theta_{q},\;\varphi_{q},\;\phi_{1}\;,\phi_{2}\;,\phi_{3}\}\;;\\ \text{(C)}:&\quad\{m_{A},\;\epsilon,\;\theta_{q},\;\varphi_{q},\;\phi_{1}\;,\phi_{2}\;,\phi_{3}\}\;.\end{split}$
We fix the values of $f_{\rm GW}$ and $m_{A}$, and scan the other model parameters in their individual parameter space in each case. The 95% exclusion limits on the $h_{0}$, $\rho_{\rm DM}$ and $\epsilon$ after integrating out all angular parameters are shown in Fig. 2. Some remarks are in order. Firstly, in the high-frequency range ($f\gtrsim 2/T_{\rm obs}$), the limits obtained using two different subtraction methods are indistinguishable, as expected, whereas linear subtraction offers better sensitivity in the low-frequency range compared to quadratic subtraction. In particular, the 95% exclusion limits are enhanced by more than one order of magnitude at $f_{\rm GW}\lesssim 6\times 10^{-10}\,{\rm Hz}$ for GWs, or at $m_{A}\lesssim 3\times 10^{-24}\,{\rm eV}$ for DPDM. Secondly, the impact of time-dependent noise from unresolved binary systems significantly decreases the sensitivity by up to an order of magnitude. Hence the inclusion of such noise is important in our analysis. In addition, this decrease is more severe in the linear subtraction scheme than in the quadratic subtraction, as we explained in Sec. II. Lastly, XG-Gaia will possess extended observation duration and improved precision. It will also have enhanced resolution capabilities for binary systems, which is crucial for suppressing the noise from unresolved binaries. It is expected that the sensitivity of XG-Gaia in detecting GWs and DPDM can be increased by more than three orders of magnitude.
V Conclusion
Astrometric measurements offer a new approach to detecting signals such as GWs and DPDM in the nHz frequency range, which are independent of PTA experiments. Subtracting stellar background motion is crucial for accurately estimating the sensitivity of such surveys. In this paper, we show that the background motion of stars can be well described by their proper motion with negligible acceleration, which motivates us to propose a linear subtraction scheme. Compared to the commonly used quadratic subtraction, linear subtraction significantly enhances the detection sensitivity in the low-frequency range. We also find that the unsolved stellar binary systems may produce large time-dependent noise, reducing the detection capability. We thus expect that the next-generation upgrade of Gaia will not only achieve higher precision but also provide better resolution of binary systems in the Milky Way.
Astrometric data can also be correlated with the current existing PTA data, allowing for a more concrete validation of the correlation pattern induced by GWs or dark matter and a more sensitive probe of macroscopic polarizations of the GWs Qin et al. (2019, 2021); Çalışkan et al. (2024). Compared to PTAs, future astrometric surveys like Roman Wang et al. (2021, 2022) is able to probe $\mu$Hz GWs or beyond and ultralight bosons with higher masses, allowing for cross-correlation with other proposed measurements in the same frequency range such as binary resonance Blas and Jenkins (2022a, b) and Doppler tracking Zwick et al. (2024), potentially closing the gap between PTA and LISA frequency ranges.
Acknowledgement
We would like to thank Huai-Ke Guo for his collaboration in the early stages of this project. We also thank Christopher J. Moore for useful email exchanges. H.A. is supported in part by the National Key R&D Program of China under Grant No. 2023YFA1607104 and 2021YFC2203100, the National Science Foundation of China (NSFC) under Grant No. 11975134, and the Tsinghua University Dushi Program. J.S. is supported by Peking University under startup Grant No. 7101302974 and the National Natural Science Foundation of China under Grants No. 12025507, No. 12150015; and is supported by the Key Research Program of Frontier Science of the Chinese Academy of Sciences (CAS) under Grants No. ZDBS-LY7003. X.W. acknowledges the Royal Society as the funding source of the Newton International Fellowship. X.X. is supported by Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC2121 “Quantum Universe” — 390833306. Y.Z. is supported by the U.S. Department of Energy under Award No. DESC0009959.
Note added: During the preparation of this paper, we became aware of the recent work by Kim (2024) and Dror and Verner (2024) which considered the astrometric effect of the ultralight scalar dark matter. Our work focuses on the detection of the ultralight vector dark matter and continuous gravitational waves, which is distinctive from their themes.
Appendix A Metric perturbation induced by DPDM
Consider a general metric perturbation $h_{\mu\nu}$, the equation of motion can be written as Carroll (2019)
$\Box h_{\mu\nu}-\partial^{\alpha}\partial_{\nu}h_{\mu\alpha}-\partial^{\alpha}\partial_{\mu}h_{\nu\alpha}+\partial_{\mu}\partial_{\nu}h+\eta_{\mu\nu}(\partial^{\alpha}\partial^{\beta}h_{\alpha\beta}-\Box h)=-16\pi GT_{\mu\nu}\;,$ (23)
where $h\equiv h_{\mu\nu}\eta^{\mu\nu}$ and $\Box\equiv\partial^{\mu}\partial_{\mu}$. We work in the Minkowski background $\eta_{\mu\nu}={\rm diag}(-1,1,1,1)$. Under the gauge fixing condition $h_{0\mu}=0$, the Einstein equations become
$-16\pi GT_{00}$ $=\partial_{0}^{2}h-(\partial_{k}\partial_{l}h_{kl}-\Box h)\;,$ (24)
$-16\pi GT_{0i}$ $=-\partial_{k}\partial_{0}h_{ik}+\partial_{i}\partial_{0}h\;,$ (25)
$-16\pi GT_{ij}$ $=\Box h_{ij}-\partial_{k}\partial_{i}h_{jk}-\partial_{k}\partial_{j}h_{ik}+\partial_{i}\partial_{j}h+\delta_{ij}\partial_{k}\partial_{l}h_{kl}-\delta_{ij}\Box h\;.$ (26)
From Eq. (26) we find
$-2\Box h+\nabla^{2}h+\partial_{l}\partial_{j}h_{jl}=-16\pi GT_{kk}\;.$ (27)
It is also convenient to derive
$-16\pi G\left(T_{ij}-\frac{\delta_{ij}}{2}T_{kk}\right)=\Box h_{ij}-\partial_{k}\partial_{i}h_{jk}-\partial_{k}\partial_{j}h_{ik}+\partial_{i}\partial_{j}h+\frac{1}{2}\delta_{ij}\partial_{k}\partial_{l}h_{kl}-\frac{1}{2}\delta_{ij}\nabla^{2}h\;.$ (28)
Next we calculate the stress-energy tensor induced by the DPDM. The virial velocity of our galaxy is $v_{\rm vir}\simeq{\cal O}(10^{-3})$, so for ultralight DPDM with mass $m_{A}\sim\mathrm{10^{-23}~{}eV}$, the de Broglie wavelength is $\lambda_{dB}\sim 4\,\mathrm{kpc}\,(10^{-23}\,\mathrm{eV}/m_{A})(10^{-3}/v_{\rm vir})$, which is of the same order as the galactic size. The coherence time is $t_{c}=4\pi/(m_{A}v^{2}_{\rm vir})=26\,\mathrm{Myr}\,(10^{-23}\,\mathrm{eV}/{m_{A}})(10^{-3}/v_{\rm vir})^{2}$. Since we are considering DPDM within a coherent time, we can write down the spatial component of dark photon field as
$\bm{A}(t,\boldsymbol{x})=A_{0}(\boldsymbol{x}){\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})e^{iEt}\right]\;,$ (29)
where $E=m_{A}\sqrt{1+\mathcal{O}(v_{\rm vir}^{2})}$ , $\boldsymbol{q}(\boldsymbol{x})$ is a complex vector that characterizes the polarization of the dark photon field, normalized as $|\boldsymbol{q}|=1$. The spatial derivative of $A(\boldsymbol{x})$ and $\boldsymbol{q}(\boldsymbol{x})$ is suppressed by virial velocity $v_{\rm vir}$. Using the equation of motion $\partial_{\mu}A^{\mu}=0$, we find that the scalar potential of the dark photon field $A^{t}$ is suppressed by $v_{\rm vir}$
$A^{t}=\mathcal{O}(v_{\rm vir})\;.$ (30)
The stress-energy tensor of vector field reads
$T_{\mu\nu}(t,\boldsymbol{x})=\eta_{\mu\nu}\left(-\frac{1}{4}F_{\rho\sigma}F^{\rho\sigma}-\frac{1}{2}m^{2}A^{\rho}A_{\rho}\right)+\eta^{\rho\sigma}F_{\mu\sigma}F_{\nu\rho}+m^{2}A_{\mu}A_{\nu}\;.$ (31)
We find that at the zeroth order of $v_{\rm vir}$, the relevant components of the stress-energy tensor are
$T_{00}^{(0)}=\rho(\boldsymbol{x})\;,\qquad T_{0k}^{(0)}=0\;,$ (32)
$T^{(0)}_{kk}=-\rho(\boldsymbol{x}){\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\,e^{2iEt}\right]\;,$
$T_{ij}^{(0)}-\frac{\delta_{ij}}{2}T_{kk}=2\rho(\boldsymbol{x}){\rm Re}\left\{\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right]e^{2iEt}\right\}\;.$
where for simplicity we defined $\rho(\boldsymbol{x})\equiv m^{2}_{A}A_{0}^{2}(\boldsymbol{x})/2$. By combining Eq. (24) and Eq. (27), we can eliminate $\partial_{l}\partial_{k}h_{kl}$ and get
$\partial_{0}^{2}h=-8\pi G\left(T_{00}+T_{kk}\right)=-8\pi G\rho(\boldsymbol{x})\left\{1-{\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\,e^{2iEt}\right]\right\}\;.$ (33)
Integrating Eq. (33), we find that the trace $h$ contains at least an oscillating component
$h^{\rm osc}=-\frac{8\pi G\rho(\boldsymbol{x})}{4m^{2}_{A}}{\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\,e^{2iEt}\right]\;,$ (34)
which suggests that $h_{ij}$ also has an oscillating part
$h_{ij}^{\rm osc}(t,\boldsymbol{x})=-h_{0}^{\rm osc}(\boldsymbol{x}){\rm Re}\left[\mathcal{H}^{\rm DP}_{ij}(\boldsymbol{x})\,e^{2iEt}\right]\;.$ (35)
Combining Eq. (28) and Eq. (32), we obtain
$-32\pi G\rho(\boldsymbol{x}){\rm Re}\left\{\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right]\,e^{2iEt}\right\}=\Box h_{ij}$ $-\partial_{k}\partial_{i}h_{jk}-\partial_{k}\partial_{j}h_{ik}+\partial_{i}\partial_{j}h$
$+\frac{1}{2}\delta_{ij}\partial_{k}\partial_{l}h_{kl}-\frac{1}{2}\delta_{ij}\nabla^{2}h\;,$ (36)
which is accurate to the zeroth order of $v_{\rm vir}$. We extract the oscillation parts on both sides of Eq. (36), which yields
$\partial_{0}^{2}{\rm Re}\left[-h_{0}^{\rm osc}(\boldsymbol{x})\mathcal{H}^{\rm DP}_{ij}(\boldsymbol{x})\,e^{2iEt}\right]=32\pi G\rho(\boldsymbol{x}){\rm Re}\left\{\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right]\,e^{2iEt}\right\}\;,$ (37)
where we neglect the terms with spatial derivatives since they are suppressed by $v_{\rm vir}$. Solving this equation, we get
$\mathcal{H}^{\rm DP}_{ij}(\boldsymbol{x})=q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\;.$ (38)
and
$h_{0}^{\rm osc}(\boldsymbol{x})=\frac{8\pi G\rho(\boldsymbol{x})}{m^{2}_{A}}\;.$ (39)
Appendix B Gauge invariance of the angular deflection
In this appendix we calculate the angular deflection of light from stars due to the oscillating metric perturbations. This part follows the discussion in Book and Flanagan (2011). In this part, we also prove the gauge invariance of the angular deflection. Consider a general metric perturbation
$\mathrm{d}s^{2}=g_{\mu\nu}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}=(\eta_{\mu\nu}+h_{\mu\nu})\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}\;.$ (40)
where we work on Minkowski background with $\eta_{\mu\nu}={\rm diag}\{-1,1,1,1\}$. We consider a photon travels from the source to the detector, then the angular deflection of light we need to calculate becomes angular difference of the observed photon momentum compared to the flat spacetime case. Without metric perturbation, we choose the detector to be at the origin $(0,\boldsymbol{0})$ and the unperturbed photon worldline to be $x_{(0)}^{\mu}=\omega_{0}(\lambda,-\lambda\boldsymbol{n})$, where $\boldsymbol{n}$ is the unit vector of the source, $\omega_{0}$ is the unperturbed photon frequency. The unperturbed photon four-momentum is $k_{(0)}^{\mu}=\omega_{0}\mathcal{N}^{\mu}$ where we define $\mathcal{N}^{\mu}=(1,-\boldsymbol{n})$. Suppose the unperturbed position of the source is $x_{s}^{i}(t)=x_{s}^{i}$, then the affine parameter of the source is $\lambda_{s}=-|\boldsymbol{x}_{s}|/\omega_{0}$.
We separate the photon worldline into unperturbed parts and the first-order perturbation
$k^{\mu}(\lambda)=k_{(0)}^{\mu}(\lambda)+k_{(1)}^{\mu}(\lambda)+\mathcal{O}(h^{2})\;,$ (41)
To the first order of $h$, the geodesic equation of photon is
$\frac{\mathrm{d}k^{\rho}}{\mathrm{d}\lambda}=-\Gamma^{\rho}_{\mu\nu}k^{\mu}_{(0)}k^{\nu}_{(0)}\;.$ (42)
We integrate from $\lambda=0$ to arbitrary $\lambda$ to get $k^{\mu}(\lambda)$
$k^{\mu}(\lambda)=k_{(0)}^{\mu}+\kappa^{\mu}-\int_{0}^{\lambda}\Gamma^{\rho}_{\mu\nu}k^{\mu}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime}\;,$ (43)
where $\kappa^{\mu}$ represents the first order perturbation term of the photon four-momentum at $\lambda=0$. Integrating Eq. (43) again gives the full trajectory of photon
$x^{\mu}(\lambda)=\lambda(k_{(0)}^{\mu}+\kappa^{\mu})-\int_{0}^{\lambda}\int_{0}^{\lambda^{\prime}}\Gamma^{\rho}_{\mu\nu}(\lambda^{\prime\prime})k^{\mu}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}+x^{\mu}(0)\;,$ (44)
where we use $x^{\mu}(0)$ to express the final position of the photon, which also refers to the position of the detector. Choosing $\lambda=\lambda_{s}$ in Eq. (44), we obtain
$\kappa^{\mu}=\frac{1}{\lambda_{s}}\left(x^{\mu}(\lambda_{s})-x^{\mu}(0)+\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}\Gamma^{\mu}_{\gamma\nu}(\lambda^{\prime\prime})k^{\gamma}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}\right)-k_{(0)}^{\mu}\;.$ (45)
To the zeroth order of perturbation, $x^{\mu}(\lambda_{s})-x^{\mu}(0)=\lambda_{s}k_{(0)}^{\mu}$, so they should cancel with each other in Eq. (45). However, in a general case with non-zero metric perturbation, we can not assume that the detector and source are static since the geodesic equations of their spatial trajectories also contain $\mathcal{O}(h)$ perturbations. Therefore we keep $x^{\mu}(\lambda_{s})$ and $x^{\mu}(0)$ in the expression of $\kappa^{\mu}$.
In addition, the null geodesic equation constraint $g_{\mu\nu}(\lambda)k^{\mu}(\lambda)k^{\nu}(\lambda)=0$ yields
$\kappa_{0}=n_{i}\kappa_{i}-\frac{1}{2\omega_{0}}h_{\mu\nu}(0)k^{\mu}_{(0)}k^{\nu}_{(0)}\;.$ (46)
To calculate the angular deflection the detector observes, we need to project the photon four-momentum into the local proper reference frame of the detector. We define the four-momentum of the detector as $e^{\nu}_{0}$ and a set of orthonormal basis vectors as $e^{\nu}_{i},i=1,2,3$. We decompose the basis vectors into unperturbed parts and the first-order perturbations
$e^{\nu}_{\alpha}=\delta^{\nu}_{\alpha}+\delta e^{\nu}_{\alpha}\;,\quad\alpha=0,1,2,3\;.$ (47)
The basis vectors of local proper reference frame satisfy
$e^{\mu}_{0}\nabla_{\mu}e^{\nu}_{\alpha}=0\;.$ (48)
Inserting Eq. (47) into Eq. (48), we find
$\partial_{0}\delta e^{\nu}_{\alpha}+\Gamma_{\alpha 0}^{\nu}=0\;.$ (49)
The orthonormal conditions of $e^{\nu}_{\alpha}$ implies
$g_{\mu\nu}e^{\nu}_{\alpha}e^{\mu}_{\beta}=\delta_{\alpha\beta}\;.$ (50)
We keep only the first order terms of Eq. (50) which gives another constraint on $\delta e^{\nu}_{\alpha}$
$h_{\mu\nu}+\eta_{\mu\gamma}\delta e^{\gamma}_{\nu}+\eta_{\nu\gamma}\delta e^{\nu}_{\mu}=0\;.$ (51)
The angular deflection of the dark photon reads
$\delta n_{i}=-\frac{g_{\nu\mu}(0)k^{\nu}(0)e^{\mu}_{i}}{\omega_{d}}-n_{i}\;,$ (52)
where $\omega_{d}=-g_{\nu\mu}(0)k^{\nu}$ is the observed photon frequency. Expanding Eq. (52) to the first order, we obtain
$\delta n_{i}=-\frac{1}{\omega_{0}}\left(h_{\nu i}(0)k^{\nu}_{(0)}+\kappa_{i}+\eta_{\nu\mu}k^{\nu}_{(0)}\delta e^{\mu}_{i}-n_{i}(h_{\nu 0}(0)k^{\nu}_{(0)}+\eta_{\nu\mu}k^{\nu}_{(0)}\delta e^{\mu}_{0}+\kappa_{0})\right)\;.$ (53)
Plugging Eq. (46) into Eq. (53), we find
$\delta n_{i}=-\frac{1}{\omega_{0}}P_{ij}\kappa_{j}-h_{\nu i}(0)\mathcal{N}^{\nu}-\mathcal{N}_{\mu}\delta e^{\mu}_{i}+n_{i}\left(h_{\nu 0}(0)\mathcal{N}^{\nu}+\mathcal{N}_{\mu}k_{d(1)}^{\mu}\right)-\frac{1}{2}n_{i}h_{\nu\mu}\mathcal{N}^{\mu}\mathcal{N}^{\nu}\;,$ (54)
where $P_{ij}=\delta_{ij}-n_{i}n_{j}$ is the projection operator. Multiplying both sides of Eq. (54) by $n^{i}$, we obtain
$n^{i}\delta n_{i}$ $=-n^{i}h_{\nu i}(0)\mathcal{N}^{\nu}-n^{i}\mathcal{N}_{\mu}\delta e^{\mu}_{i}+h_{\nu 0}(0)\mathcal{N}^{\nu}+\mathcal{N}_{\mu}\delta e_{0}^{\mu}-\frac{1}{2}h_{\nu\mu}\mathcal{N}^{\mu}\mathcal{N}^{\nu}$
$=-n^{i}\mathcal{N}_{\mu}\delta e^{\mu}_{i}+\mathcal{N}_{\mu}\delta e_{0}^{\mu}+\frac{1}{2}h_{\nu\mu}\mathcal{N}^{\mu}\mathcal{N}^{\nu}$
$=\mathcal{N}_{\mu}\mathcal{N}^{\nu}\delta e^{\mu}_{\nu}+\frac{1}{2}h_{\nu\mu}\mathcal{N}^{\mu}\mathcal{N}^{\nu}$
$=0\;.$ (55)
Note that in the last step we use Eq. (51). Now $n^{i}\delta n_{i}=0$ shows that $\delta n_{i}$ is orthogonal to $n^{i}$, so we can multiply $P_{ij}$ to each term in the expression of $\delta n_{i}$, which reads
$\delta n_{i}=-P_{ij}\left(\frac{1}{\omega_{0}}\kappa_{j}+h_{\nu j}(0)\mathcal{N}^{\nu}+\mathcal{N}_{\mu}\delta e^{\mu}_{j}\right)\;.$ (56)
Then we prove that this expression is gauge-invariant. Consider a transformation of coordinates $x^{\prime\mu}=x^{\mu}+\xi^{\mu}(x)$, it induces a gauge transformation of linearized theory
$h_{\mu\nu}(x)\rightarrow h_{\mu\nu}(x)-(\partial_{\mu}\xi_{\nu}+\partial_{\nu}\xi_{\mu})\;.$ (57)
We check the gauge transformation of each term in $\delta n_{i}$. We use $\Delta$ to represent the change of each term after gauge transformation. Notice that $\Delta\Gamma^{\rho}_{\mu\nu}=-\partial_{\mu}\partial_{\nu}\xi^{\rho}$, $\Delta x^{\mu}(\lambda)=\xi^{\mu}(\lambda)$, we have
$\Delta\kappa^{\mu}$ $=\frac{1}{\lambda_{s}}\left(\xi^{\mu}(\lambda_{s})-\xi^{\mu}(0)-\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}\partial_{\gamma}\partial_{\nu}\xi^{\mu}(\lambda^{\prime\prime})k^{\gamma}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}\right)$
$=k^{\nu}_{(0)}\partial_{\nu}\xi^{\mu}(0)=\omega_{0}\mathcal{N}^{\nu}\partial_{\nu}\xi^{\mu}(0)\;.$ (58)
In the rest of the discussion we ignore the coordinate dependence of $\xi^{\mu}$ since now all the $\xi^{\mu}(\lambda)$ are defined on $\lambda=0$. Here we use $k^{\nu}_{(0)}\partial_{\nu}\xi^{\mu}=\mathrm{d}\xi^{\mu}/\mathrm{d}\lambda$ to simplify the integration.
From Eq. (49) we find
$\partial_{0}(\Delta\delta e^{\nu}_{i})=-\Delta\Gamma_{i0}^{\nu}=\partial_{0}\partial_{i}\xi^{\nu}\;,$ (59)
which means $\Delta\delta e^{\nu}_{i}=\partial_{i}\xi^{\nu}+f$. $f$ is an arbitrary function that has no dependence on time. We ignore $f$ because it is not correlated with gauge transformation. Insert these expressions into $\delta n_{i}$, we can get the gauge transformation of the angular deflection.
$\Delta\delta n_{i}=-P_{ij}\left(\mathcal{N}^{\nu}\partial_{\nu}\xi_{j}-(\partial_{j}\xi_{\nu}+\partial_{\nu}\xi_{j})\mathcal{N}^{\nu}+\mathcal{N}_{\mu}\partial_{j}\xi^{\mu}\right)=0\;.$ (60)
Therefore the angular deflection $\delta n_{i}$ is gauge-invariant. Now we choose a typical gauge to calculate $\delta n_{i}$. We use synchronous gauge which is defined by $h_{00}=h_{0i}=0$. The geodesic equations of the four-momentum of the detector is reduced to
$\partial_{0}\delta e^{\nu}_{\mu}=-\Gamma_{\mu 0}^{\nu}=-\frac{1}{2}\partial_{0}h_{\mu}^{\nu}\;,$ (61)
which gives
$\delta e^{\nu}_{\mu}=-\frac{1}{2}h_{\mu}^{\nu}\;.$ (62)
For $\mu=0$, $\delta e^{\nu}_{0}=0$. Thus the spatial components of the four-momentum of the detector do not contain first order perturbations. Then we can treat the detector as a static object and $x^{\mu}(0)$ becomes a constant. Similarly, $x^{\mu}(\lambda_{s})$ is also a constant, and they satisfy $x^{\mu}(\lambda_{s})-x^{\mu}(0)=\lambda_{s}k^{\mu}_{(0)}$. The final expression of angular deflection is
$\delta n_{i}$ $=-P_{ij}\left(\frac{\eta_{j\mu}}{\omega_{0}\lambda_{s}}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}\Gamma^{\mu}_{\gamma\nu}(\lambda^{\prime\prime})k^{\gamma}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}-\frac{1}{2}h_{jk}(0)n^{k}\right)$
$=-P_{ij}\left(\frac{\omega_{0}}{\lambda_{s}}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}\left(-n^{l}\partial_{0}h_{jl}+n^{l}n^{k}\partial_{k}h_{jl}-\frac{1}{2}n^{l}n^{k}\partial_{j}h_{kl}\right)\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}-\frac{1}{2}h_{jk}(0)n^{k}\right)\;.$ (63)
Using the identity
$\frac{\mathrm{d}}{\mathrm{d}\lambda}h_{jl}=\omega_{0}\left(\partial_{0}h_{jl}-n^{k}\partial_{k}h_{jl}\right)\;,$ (64)
we can integrate the first two terms in Eq. (63)
$\delta n_{i}$ $=-P_{ij}\left(\frac{1}{\lambda_{s}}\int_{0}^{\lambda_{s}}\left(-n^{l}h_{jl}(\lambda^{\prime})+n^{l}h_{jl}(0)\right)\mathrm{d}\lambda^{\prime}-\frac{\omega_{0}}{2\lambda_{s}}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}n^{l}n^{k}\partial_{j}h_{kl}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}-\frac{1}{2}h_{jk}(0)n^{k}\right)$
$=-P_{ij}\left(-\frac{1}{\lambda_{s}}\int_{0}^{\lambda_{s}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}-\frac{\omega_{0}}{2\lambda_{s}}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}n^{l}n^{k}\partial_{j}h_{kl}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}+\frac{1}{2}h_{jk}(0)n^{k}\right)\;.$ (65)
The angular deflection is separated into three terms
$\delta n_{1i}$ $=\frac{1}{\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}\;,$ (66)
$\delta n_{2i}$ $=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}\;,$ (67)
$\delta n_{3i}$ $=-\frac{1}{2}P_{ij}h_{jk}(0)n^{k}\;.$ (68)
Now we estimate the relative size of these three terms. We introduce the coherent length $l_{c}$ which is $l_{c}=2\pi/(m_{A}v_{\rm vir})$. Without loss of generality, we consider the metric perturbation that takes the following form
$h_{ij}(x,t)=c\mathcal{H}_{ij}\rho(\boldsymbol{x})\cos(2m_{A}t+2\alpha)\;.$ (69)
where $\mathcal{H}_{ij}$ is an order one tensor, $\rho(\boldsymbol{x})$ refers to the local dark matter energy density, $c=-8\pi G/m_{A}^{2}$. For ultralight dark matter, a core structure can form at the center of the Galaxy with a radius of $\sim l_{c}$. The energy density of the core can be much larger than the energy density outside the core. We expect that the photon emitted by a star inside the core may have larger $\delta n_{1i}$ and $\delta n_{2i}$ since part of the worldline of the photon is inside the core. However, we will see in the following calculation that even the star is inside the core, $\delta n_{3i}$ still dominates $\delta n_{i}$.
Consider the case that the source of photon is inside the core. We suppose that at $\lambda=\lambda_{b}$, the photon touches the boundary of the core. Therefore we expect $\omega_{0}|\lambda_{s}-\lambda_{b}|\sim l_{c}$. First we calculate $\delta n_{1i}$. We separate $\delta n_{1i}$ into two parts
$\delta n_{1i}=\frac{c}{\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}+\frac{c}{\lambda_{s}}P_{ij}\int_{0}^{\lambda_{b}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}\;,$ (70)
Since the $h_{ij}$ inside the core is much larger, we only calculate the first integration. In addition, the integral length is approximately one coherent length, so we can use Eq. (69) to express $h_{ij}$. Insert Eq. (69) into $\delta n_{1i}$, we obtain
$\delta n_{1i}=\frac{c}{\lambda_{s}}P_{ij}n^{l}\mathcal{H}_{jl}\int_{\lambda_{b}}^{\lambda_{s}}\rho(\lambda^{\prime})\cos(2m_{A}\omega_{0}\lambda^{\prime}+2\alpha)\mathrm{d}\lambda^{\prime}\sim P_{ij}n^{l}\mathcal{H}_{jl}\frac{c\rho_{c}}{m_{A}\omega_{0}\lambda_{s}}\;,$ (71)
where $\rho_{c}$ refers to the energy density of the soliton core. For simplicity, we use NFW profile to estimate $\rho_{c}$. Suppose the energy density of DPDM near the Earth is $\rho_{0}$. The NFW profile takes the form as McMillan (2011)
$\rho_{h}=\frac{\rho_{h,0}}{x(1+x)^{2}}\;,$ (72)
where $x=r/r_{h}$ and $r_{h}=17\mathrm{kpc}$. We use the energy density near $r=l_{c}$ to estimate the energy density of the core. Then the ratio of $\rho_{c}$ and $\rho_{0}$ is
$\frac{\rho_{c}}{\rho_{0}}\sim\frac{l_{e}}{l_{c}}\sim l_{e}m_{A}v_{\rm vir}\;,$ (73)
where $l_{e}=8\mathrm{kpc}$ is the distance between the Earth and the center of the Milky Way. Using the expression of $\rho_{c}$, we obtain
$\delta n_{1i}\sim P_{ij}n^{l}\mathcal{H}_{jl}\frac{c\rho_{0}v_{\rm vir}l_{e}}{x_{s}}\sim cP_{ij}n^{l}\mathcal{H}_{jl}v_{\rm vir}\rho_{0}\sim v_{\rm vir}\delta n_{3i}\;,$ (74)
$x_{s}$ is the distance from the star to the observer. $l_{e}/x_{s}$ is an order one parameter, so we ignore it. The virial velocity is $v_{\rm vir}\sim 10^{-3}$. Therefore, $\delta n_{1i}$ is suppressed compared to $\delta n_{3i}$.
Next we consider $\delta n_{2i}$. We exchange the order of integration in $\delta n_{2i}$, which yields
$\delta n_{2i}$ $=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}\int_{\lambda^{\prime\prime}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime}n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})$
$=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})$
$=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})+\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{b}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})\;.$ (75)
Again we only calculate the first integral. Inside the core, we find $\omega_{0}|\lambda_{s}-\lambda^{\prime\prime}|\sim l_{c}=\frac{1}{m_{A}v_{\rm vir}}$ and $\partial_{i}h_{jk}\sim m_{A}v_{\rm vir}h_{jk}$. Then we have
$\delta n_{2i}$ $=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})$
$\sim\frac{1}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}\left(\frac{1}{m_{A}v_{\rm vir}}\right)n^{l}n^{k}m_{A}vh_{kl}(\lambda^{\prime\prime})$
$\sim\frac{1}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}n^{l}n^{k}h_{kl}(\lambda^{\prime\prime})\sim\delta n_{1i}\;,$ (76)
where we find $\omega_{0}(\lambda_{s}-\lambda^{\prime\prime})$ and the spatial derivative of $h_{ij}$ cancel with each other, making $\delta n_{1i}$ and $\delta n_{2i}$ at the same order. Then similar to $\delta n_{1i}$, $\delta n_{2i}$ is also suppressed compared to $\delta n_{3i}$.